OpenStudy (anonymous):

solve by completing the square 6x^2+19x=7

OpenStudy (anonymous):

6x^2+19x=7 use quadratic formula

OpenStudy (anonymous):

sorry i didn read ur question completin the square.....mmmmm....lemi see... al have 2 refer for the answer

OpenStudy (anonymous):

To complete the square you have something like the following: \[6x^2+19x=7 \implies 6(x^2 + \frac{19}{6}x) = 7\] Then you take half the x term coefficient and add it to both sides.

OpenStudy (anonymous):

\[x^2 + \frac{19}{6}x +\frac{19^2}{12^2} = \frac{7}{6} + \frac{19^2}{12^2}\] \[\implies (x + \frac{19}{12})^2 = \frac{529}{144}\]

OpenStudy (anonymous):

Kind of a nasty quadratic to be doing this with though

OpenStudy (anonymous):

so wats the answer?

OpenStudy (anonymous):

yeah its a crappy question , I have a feeling its not solution its so crazy